# EgoBlind: Towards Egocentric Visual Assistance for the Blind

### v2

@@ 3.2 Statistic Analysis
Questions. Figure 3(a) shows that more than 80% of the questions are short in 10 words, with an average length of 7.9. The relatively shorter questions could be attributed to the use of more spoken language in practice. The word cloud in Figure 3(e) shows that the questions are from blind users’ first-person perspective (“I”, “me”, “my”) and feature referential concepts, such as “this”, “there”, “that”, “it” and “now”, which often demand on engaging in the scenarios, capturing the questioner’s hand motion and eye gaze, and reasoning the spatio-temporal contexts for understanding the questions. Figure 3(c) shows that blind people are often interested in locating something (“where is”) and checking for potential safety issues (“are there any obstacles”).
@@
Answers. Figure 3(b) shows that the answers are relatively longer with an average of 5.7 words. Near half of the answers have more than 3 words, with 19.4% over 10 words. A detailed study shows that the answers for tool use are significantly longer and more complex than other assistive questions; the average answer length is 12.1 for tool use while 4 to 6 for others. The most frequent answers are “yes” (13.0%), followed by “I do not know” (10.7%) and “no” (9.6%). The answer of “I do not know” is due to poor visual quality or the answer not framed by the blind users. Additionally, it may be attributed to our real-time QA task setting, i.e., the answer is not visible up to the question moment. We keep these unanswerable questions to evaluate if models can reject to answer rather than hallucinate potentially malicious answers. For better analysis, we omit these answers and analyze the remaining frequent answers and answer word clouds in Figure 3(d) and 3(f), respectively. We find that the answers are mostly about directions, navigation and locations. It is worth mentioning that the locations and directions are often relative to the camera wearers (questioners).
@@
Videos. Figure 3(g) shows that the vast majority of the videos (82.3%) are within 1 minute, with 17.7% exceeding 1 minute and 7.1% running longer than 2 minutes. The mean video duration is 40.0 seconds, reflecting the average duration each time a visually-impaired person recording their daily moment. Additionally, we conduct a comparison between the egocentric video content captured by sighted people (e.g., Ego4D [32] videos) and those are blind (examples are presented in Appendix A.3). We find that the differences are from the following major aspects: 1) Composition and Focus: Blind individuals often produce footage where subjects are off-center or out of focus (refer to Figure 1 and Appendix). 2) Camera Orientation and Stability: Blind individuals hard to maintain consistent camera orientation, potentially leading to tilted or unstable footage. 3) Environmental Awareness: Blind individuals inadvertently capture obstructed views or poorly lit scenes. Nevertheless, the video resolution is higher, possibly because the videos are crawled from video sharing platforms.
3.3 Dataset Comparison
Unlike most Ego-VQA datasets (the first block of Table 3), which focus on general-purpose visual understanding, EgoBlind emphasizes assisting blind individuals in different assistance aspects. Compared to other assist-oriented datasets (the middle block), EgoBlind covers the wide aspects of visual assistance for the blind people, versus episodic memory in QAEgo4D [35] or tool use in AssistQ [31] or scene-text QA in EgoTextVQA [38].
Compared to the blind-related QA datasets (the bottom block), EgoBlind advances VizWiz [24] by enabling egocentric live QA with more visual information about real-world dynamics and episodic memory support. Although the videos in the VIEW-QA [13] dataset offer 360-degree visual information, the videos and questions are simulated by singted people mimicking blind individuals, thus lacking sufficient practical characteristics for the blind community, especially for those with long blind age. In contrast, the videos in EgoBlind are entirely filmed by blind individuals themselves, authentically reflecting their real-life scenarios. Moreover, the questions in EgoBlind are posed or verified by blind individuals of different blind ages, further ensuring the dataset’s practicality and relevance. Additionally, EgoBlind supports a live QA setting with timestamp annotation of each question and the answers are based on the video contents prior to the question timestamp.
@@ 3.3 Dataset Comparison
Table 3: Dataset Comparison. BV: Video captured by blind people. BQ: Question posed or verified by blind people. RT: Answe based on the video content up to the question timestamp. MA: Multiple reference answers for each question. QC: Question Category. OE/MC: Open-ended/Multi-choice.
@@ 4 Experiment
4.1 Experimental Setup
Evaluation. We used GPT-4o mini as a validation tool to assess the model performances on EgoBlind test set. It evaluates the semantic similarity between the models’ predictions and the ground truth (reference answers). Building on previous work, we established two evaluation metrics for the models’ predictions: accuracy (0-100%, the percentage of ‘yes’ answers evaluated) and score (0-5, degree of similarity to the matching results). Specially, for each question, a correct prediction is identified if it achieves yes-response with any one of the reference answers. Note that the evaluation prompts are manually finetuned to reach a maximal score agreement (0.88) between human and AI reviewers (details in Appendix B.1).
Model Setup To benchmark the challenges carried by EgoBlind, we comprehensively analyze 15 contemporary MLLMs, including 12 open-source models and 3 closed-source models (via APIs). The choice of the open-source models are based on the following criteria: 1) Achieve SOTA results on common video QA benchmarks such as NExT-QA [8], VideoChatGPT-Bench [18] and Video-MME [19]. Corresponding models are LLaVA-OV [15], CogVLM2-Video [47], Video-LLaMA3 [16], InternVL2.5 [14]; 2) Achieve SOTA results on general-purpose egocentric VideoQA benchmarks such as EgoSchema [11], VidEgoThink [10] and EgoMemoria [20]. Corresponding models are LLaVA-Video [48], MiniCPM-V 2.6 [49] and Qwen2.5-VL[4]; 3) Achieve SOTA on the image blind QA dataset VizWiz [24], such as Video-LLaVA [18], LLaMA-VID[22] and VILA 1.5 [23]. Noteworthy, for each question, we uniformly sample the video content up to the question timestamp for answer prediction, thus to match the live QA task setting. Moreover, we design customized prompts tailored for answering blind users’ questions for better performance. Additionally, we obtain human performance by inviting 3 university students who did not participant in annotation.
Table 4: Model behaviors (accuracy / score) on EgoBlind.
Methods LLM Res. #F Tool Use Information Navigation Safety Communication Resource Average
Human - - - 70.4 / 3.4 87.0 / 4.3 83.1 / 3.9 91.9 / 4.5 94.7 / 4.7 96.6 / 4.6 87.4 / 4.2
@@ 4.1 Experimental Setup
ShareGPT4Video [46] LLaMA3-8B - 2fps 36.0 / 2.3 34.5 / 2.0 18.6 / 1.5 43.5 / 2.3 40.3 / 2.2 26.0 / 1.7 33.9 / 2.0
CogVLM2-Video [47] LLaMA3-8B 2242 24 37.4 / 2.4 45.8 / 2.4 15.5 / 1.4 51.4 / 2.6 44.4 / 2.4 34.1 / 2.0 41.2 / 2.3
Video-LLaMA3 [16] Qwen2.5-7B ori 1fps 46.9 / 2.5 54.5 / 2.8 38.9 / 2.2 53.6 / 2.8 48.6 / 2.6 53.8 / 2.8 51.3 / 2.7
LLaVA-OV [15] Qwen2-7B 3842 16 59.9 / 3.1 57.1 / 3.0 31.4 / 2.1 65.9 / 3.3 58.3 / 3.0 50.2 / 2.7 55.1 / 2.9
InternVL2.5-8B [14] InternLM2_5-Chat-7B 4482 8 59.2 / 3.2 57.9 / 3.0 43.8 / 2.6 60.1 / 3.0 44.4 / 2.4 54.3 / 2.9 55.8 / 2.9
InternVL2.5-26B [14] InternLM2_5-Chat-20B 4482 8 69.4 / 3.4 58.6 / 3.0 49.5 / 2.7 57.0 / 2.9 43.1 / 2.3 53.2 / 2.8 56.6 / 2.9
MiniCPM-V 2.6 [49] Qwen2-7B 3842 1fps 51.0 / 2.5 49.8 / 2.6 38.4 / 2.2 34.2 / 1.9 37.5 / 2.0 42.8 / 2.2 43.6 / 2.3
Qwen2.5-VL [4] Qwen2.5-7B ori 1fps 56.5 / 3.1 53.8 / 2.8 31.2 / 2.1 53.5/ 2.8 41.7 / 2.4 38.2 / 2.2 49.0 / 2.7
LLaVA-Video [48] Qwen2-7B 3842 1fps 39.5 / 2.2 54.7 / 2.8 34.8 / 2.1 61.9 / 3.1 48.6 / 2.6 49.7 / 2.6 52.0 / 2.7
Video-LLaVA [18] Vicuna-7B 2242 8 21.8 / 1.8 42.4 / 2.3 21.4 / 1.5 48.5 / 2.5 41.7 / 2.2 36.4 / 2.0 39.1 / 2.2
LLaMA-VID [22] Vicuna-7B 2242 1fps 33.3 / 2.3 41.8 / 2.3 19.6 / 1.5 52.2 / 2.7 40.3 / 2.2 37.0 / 2.1 40.1 / 2.2
VILA1.5 [23] LLaMA3-8B 3362 8 45.6 / 2.6 49.6 / 2.6 25.3 / 1.6 58.5 / 2.9 50.0 / 2.6 41.0 / 2.3 47.2 / 2.5
@@
Gemini 1.5 Flash - ori - 74.2 / 3.7 59.4 / 3.0 44.3 / 2.6 60.4 / 3.0 58.3 / 2.9 54.3 / 2.9 57.9 / 3.0
Gemini 2.0 Flash - ori - 57.8 / 3.0 58.8 / 3.0 48.4 / 2.7 56.1 / 2.8 52.8 / 2.7 48.6 / 2.6 55.7 / 2.9
Gemini 2.5 Flash - ori - 66.7 / 3.1 57.9 / 3.0 50.3 / 2.7 55.1 / 2.8 48.6 / 2.5 58.4 / 3.0 56.3 / 2.9
GPT-4o - ori 1fps 65.3 / 3.4 64.7 / 3.3 51.0 / 2.9 58.8 / 3.0 62.4 / 3.1 55.5 / 2.9 60.6 / 3.1
@@ 4.2 Benchmarking Analysis
4.2.1 Overall Observations
Table 4 presents the performances of different models. We summarize the following observations:
@@ 4.2.1 Overall Observations
  • The models that are superior at general-purpose egocentric VQA (e.g., LLaVA-Video) and image blind-VQA (e.g., VILA1.5) are not the best-performing, depicting unique challenges of EgoBlind.
  • No single model wins across all question types. Answering “Navigation” questions is the most challenging task for almost all models, indicating a significant limitation of MLLMs in this field.
  • While most other models struggle in answering questions about tool use, Gemini 1.5 even surpasses human performance, demonstrating its rich knowledge outweighing human individuals.
  • Stronger LLMs and larger visual resolution often bring better performance, while more frames do not always help (e.g., 8 frames are enough for InternVL to surpass other open-source models.).
4.2.2 Specific Challenges
@@
User Intention. Reasoning user intentions behind the questions is key for effective assistance. Yet, all models fall short in such capability. For example, in Figure 4(a) and (c), all models fail to generate helpful answers for the blind, though the answers are objectively correct to the visual contents. It is worth mentioning that we have explicitly prompted the models to answer questions of blind individuals to provide visual assistance.
Spatio-temporal Orientation. MLLMs exhibit significant shortcomings in weaving together the temporal frames to reason the spatial orientations relative to the users’ real-time location. A typical example is shown in Figure 4(b), where indoor navigation is requested by the blind user. The escalator was framed prior to the question moment, and the models have to retrieve the related moment and reason the users’ orientation change after that moment, which shows extreme challenge to all MLLMs.
@@
We also analyze other challenges related to streaming VQA, scene text recognition, and referential words in spoken languages in the Appendix B.2.
@@
InternVL2.5-26B 56.6 55.5 $\downarrow$ 1.1
VILA1.5 47.2 45.8 $\downarrow$ 1.4
GPT-4o 60.6 60.2 $\downarrow$ 0.4
@@
InternVL2.5-26B 56.6 54.7 $\downarrow$ 1.9
Gemini 1.5 Flash 57.9 52.5 $\downarrow$ 5.4
GPT-4o 60.6 56.8 $\downarrow$ 3.8
@@
Qwen2.5-VL 49.0 53.5 $\uparrow$ 4.5
LLaVA-OV 55.1 57.5 $\uparrow$ 2.4
InternVL2.5-8B 55.8 62.1 $\uparrow$ 6.3
@@ 4.3 Other Investigations
Our investigations aim to answer three questions: 1) Is a single frame at the question timestamp enough to answer the question? 3) Whether it is helpful to explicitly prompt the models that the questions are asked by blind users for visual assistance? 2) Whether using our training data for instruction tuning can improve the performance?
Table 4.2.2 shows that replacing multi-frame inputs with a single frame at the question timestamp degenerates model performances, though a single frame brings competitive results (which is reasonable as people often ask about their current visual environment.). Table 4.2.2 shows that substituting the blind-specific prompt with a normal VQA prompt significantly jeopardize the model performance. This reveals that answering questions from blind people sparks new challenges compared with QA about general visual understanding. Table 4.2.2 shows that using our training data for instruction tuning (LoRA finetune) can remarkably improve the model performances, demonstrating the importance of our training data towards improvements. We present other details in Appendix B.3.

### v3

@@ 3.4 Statistic Analysis
Questions. Figure 3(a) shows that more than 80% of the questions are shorter than 10 words, with an average length of 7.9 words. The relatively short questions could be attributed to the use of more spoken language in practice. The word cloud in Figure 3(e) shows that the questions are from blind users’ first-person perspective (“I”, “me”, “my”) and feature referential concepts, such as “this”, “there”, “that”, “it” and “now”, which requires context-specific interpretation. Figure 3(c) shows that blind people are often interested in locating something (“where is”) and checking for potential safety issues (“are there any obstacles”).
@@
Answers. Figure 3(b) shows that the answers have an average length of 5.7 words. Near half of the answers have more than 3 words, with 19.4% over 10 words. A detailed study shows that the answers for tool use are significantly longer and more complex than other assistive questions; the average answer length is 12.1 words for tool use and 4 to 6 words for others. The most frequent answers are “yes” (13.0%), followed by “I do not know” (10.7%) and “no” (9.6%). The answer of “I do not know” is due to poor visual quality or the answer not framed by the blind users. Additionally, it may be attributed to our online QA task setting, i.e., the answer is not visible up to the question moment. We keep these unanswerable questions to evaluate if models can reject to answer rather than hallucinate potentially malicious answers. For better analysis, we omit these answers and analyze the remaining frequent answers and answer word clouds in Figure 3(d) and 3(f), respectively. We find that the answers are mostly about directions, navigation and locations. It is worth mentioning that the locations and directions are often relative to the camera wearers (questioners).
@@
Videos. Figure 3(g) shows that the vast majority of the videos (82.3%) are within 1 minute, with 17.7% exceeding 1 minute and 7.1% running longer than 2 minutes. The mean video duration is 40.0 seconds, reflecting the average duration each time a visually-impaired person recording their daily moment. Additionally, we conduct a comparison between the egocentric video content captured by sighted people (e.g., Ego4D [35] videos) and those are blind (examples are presented in Appendix A.3). Differences arise from 1) Composition and Focus: Blind individuals often produce footage where subjects are off-center or out of focus (refer to Figure 1 and Appendix A.3). 2) Camera Orientation and Stability: Blind individuals struggle to maintain consistent camera orientation, leading to potentially tilted or unstable footage. 3) Environmental Awareness: Blind individuals inadvertently capture obstructed views or poorly lit scenes. Nevertheless, the video resolution is higher, possibly because the videos are crawled from content creators on video sharing platforms.
@@
Table 3: Ego-VQA dataset comparison. BV: Video captured by the blind. BQ: Question posed or verified by the blind. RT: Answer based on the video content up to the question timestamp. MA: Multiple answers for each question. QC: Question Category. OE/MC: Open-Ended/Multi-Choice.
@@
3.5 Dataset Comparison
Unlike most Ego-VQA datasets (the first block of Table 3) which focus on general-purpose visual understanding, EgoBlind emphasizes assisting blind individuals. Compared to other assist-oriented datasets (the middle block), EgoBlind covers the wide aspects of visual assistance for the blind in daily life, versus episodic memory in QAEgo4D [38], tool use in AssistQ [34], manipulation in HoloAssist [51] (proactive assistance), or scene-text reading in EgoTextVQA [41].
Compared to the blind-related QA datasets (the bottom block), EgoBlind advances VizWiz [15] by enabling egocentric live QA with more visual information about real-world dynamics and episodic memory support. Although the videos in the VIEW-QA [17] dataset offer 360-degree visual information, the videos and questions are simulated by seeing actors. In contrast, the videos in EgoBlind are entirely filmed by blind or visually impaired individuals themselves, authentically reflecting their real-life scenarios. Moreover, the questions in EgoBlind are posed or verified by blind individuals of different blind ages, further ensuring the dataset’s practicality and relevance. Finally, EgoBlind supports an online QA setting with timestamp annotations for each question; answers are based on the video contents prior to the question timestamp.
@@ 4 Experiment
Evaluation. Following popular practices for LLMs [21, 52], we use the GPT score as a metric for evaluating generated answers. Specifically, we prompt GPT-4o mini to assess the semantic similarity between a models’ prediction and the ground truth (reference answers), and answer with ’yes’ if they are judged as the same. We then obtain the accuracy (0-100%) as the percentage of ‘yes’ answers evaluated. Noteworthy, for each question, a correct prediction is identified if it achieves yes-response with any one of the reference answers. Moreover, the evaluation prompts are manually finetuned to reach a maximal agreement (0.88) between human and AI reviewers (details in Appendix B.1).
Model Setup. To benchmark the challenges carried by EgoBlind, we comprehensively analyze 16 contemporary MLLMs, including 12 open-source models and 4 closed-source ones (via APIs). The choice of the open-source models are based on the following criteria: 1) Achieve SOTA results on common video QA benchmarks such as NExT-QA [8], VideoChatGPT-Bench [21] and Video-MME [22]. Corresponding models are LLaVA-OV [53], CogVLM2-Video [54], Video-LLaMA3 [48], InternVL2.5 [18]; 2) Achieve SOTA results on general-purpose egocentric VideoQA benchmarks such as EgoSchema [11], VidEgoThink [10] and EgoMemoria [23]. Corresponding models are LLaVA-Video [55], MiniCPM-V 2.6 [56] and Qwen2.5-VL[4]; 3) Achieve SOTA on the image blind QA dataset VizWiz [15], such as Video-LLaVA [21], LLaMA-VID[25] and VILA 1.5 [26]. Notably, for each question, we uniformly sample the video content up to the question timestamp for answer prediction, thus to match the live QA task setting. Moreover, we design customized prompts tailored for answering blind users’ questions for better performance. Additionally, we obtain human performance by inviting 3 university students who did not participant in annotation.
Table 4: QA accuracy of different models on EgoBlind.
Methods LLM Size #F Tool Info. Navi. Safe Com. Res. Overall
Human - - - 70.4 87.0 83.1 91.9 94.7 96.6 87.4
@@
ShareGPT4Video [50] LLaMA3-8B ori 16 25.5 32.6 20.7 43.3 38.9 28.3 32.9
CogVLM2-Video [54] LLaMA3-8B 2242 24 32.2 44.5 14.0 52.7 43.1 32.4 40.3
Video-LLaMA3 [48] Qwen2.5-7B ori 1fps 53.0 51.9 38.1 50.6 41.7 50.3 49.2
InternVL2.5-8B [18] InternLM2_5-7B 4482 8 61.1 54.6 42.2 58.0 44.4 52.6 53.5
LLaVA-OV [53] Qwen2-7B 3842 16 61.1 56.4 29.5 65.8 58.3 50.9 54.5
InternVL2.5-26B [18] InternLM2_5-20B 4482 8 72.5 56.9 47.4 54.1 43.1 53.2 55.0
MiniCPM-V 2.6 [56] Qwen2-7B 3842 1fps 53.7 46.5 37.8 28.9 37.5 41.0 40.7
Qwen2.5-VL [4] Qwen2.5-7B ori 1fps 51.0 50.1 28.2 48.5 43.1 38.2 45.5
LLaVA-Video [55] Qwen2-7B 3842 1fps 44.3 53.4 32.6 62.0 50.0 49.7 51.5
Video-LLaVA [21] Vicuna-7B 2242 8 22.8 41.2 21.2 47.2 38.9 35.3 38.1
LLaMA-VID [25] Vicuna-7B 2242 1fps 32.2 40.5 20.7 49.4 36.1 41.6 39.1
VILA1.5 [26] LLaMA3-8B 3362 8 49.7 50.5 25.9 60.6 47.2 41.0 48.2
@@
Gemini 2.0 Flash - ori 32 61.1 54.5 50.5 39.1 47.2 49.1 49.9
Gemini 1.5 Flash - ori 32 72.5 54.4 43.5 50.6 38.9 45.7 51.8
Gemini 2.5 Flash - ori 32 67.1 57.6 47.7 57.8 47.2 50.3 56.0
GPT-4o - ori 32 66.4 61.2 52.6 58.8 47.2 62.4 59.3
@@ 4.1 Benchmarking Analysis
(2) The models that are superior at general-purpose egocentric VQA (e.g., LLaVA-Video) and image blind-VQA (e.g., VILA1.5) are not the best-performing, depicting unique challenges of EgoBlind.
(3) No single model wins across all question types. Answering “Navigation” questions is the most challenging task for almost all models, indicating a significant limitation of MLLMs in this field.
(4) While most other models struggle in answering questions about tool use, Gemini 1.5 even surpasses human performance, demonstrating its rich knowledge outweighing human individuals.
(5) Stronger LLMs and larger visual resolution often bring better performance, while more frames do not always help (e.g., 8 frames are enough for InternVL to surpass other open-source models.).
4.2 Assistance-Related Challenges
@@
User Intention. Reasoning about user intentions behind the questions is key for effective assistance. Yet, all models fall short in such a capability. For example, in Figure 4(a) and (c), all models fail to generate helpful answers for the blind, though the answers are objectively correct to the visual contents. It is worth mentioning that this is despite explicitly prompting the models to answer questions of blind individuals to provide visual assistance.
Spatial Orientation Change. MLLMs exhibit significant shortcomings in weaving together the temporal frames to reason the spatial orientations relative to the users’ real-time location. A typical example is shown in Figure 4(b), where indoor navigation is requested by the blind user. The escalator was framed prior to the question moment, and the models have to retrieve the related moment and reason the users’ orientation change after that moment, which shows extreme challenge to all MLLMs.
@@
We also analyze other challenges related to streaming VQA, scene text recognition, and referential words in spoken languages in Appendix B.2.
@@
InternVL2.5-26B 55.0 53.4 $\downarrow$ 1.6
VILA1.5 48.2 47.2 $\downarrow$ 1.0
GPT-4o 59.3 58.2 $\downarrow$ 1.1
@@
InternVL2.5-26B 55.0 54.0 $\downarrow$ 1.0
Gemini 2.5 Flash 56.0 55.2 $\downarrow$ 0.8
GPT-4o 59.3 58.1 $\downarrow$ 1.2
@@
Qwen2.5-VL 45.5 50.2 $\uparrow$ 4.7
LLaVA-OV 54.5 57.4 $\uparrow$ 2.9
InternVL2.5-8B 53.5 58.1 $\uparrow$ 4.6
(a) InternVL2.5-26B.
(b) Gemini 2.5 Flash.
(c) GPT-4o.
Figure 5: Uniform sampling vs. single frame (Q-Frame) input at the question timestamp. The overall QA accuracy declines slightly when replacing video with a Q-Frame.
(a) InternVL2.5-26B.
(b) Gemini 2.5 Flash.
(c) GPT-4o.
Figure 6: Blind aware prompt (Appendix Table 12) vs. Normal VQA prompt (Appendix Table 11). The overall accuracy declines without blind-specific prompting.
(a) Qwen2.5-VL.
(b) LLaVA-OV.
(c) InternVL-2.5-8B.
Figure 7: Simple finetuning with EgoBlind training data can notably boost QA performances, but it is ill-suited for answering “tool-use” questions.
@@ 4.3 Other Investigations
Our investigations aim to answer three questions: 1) Is a single frame at the question timestamp sufficient to answer the question? 3) Does explicitly prompting models to assist blind users benefit performance? 2) Can instruction tuning with our training data improve performance?
Table 4.2 shows that replacing multi-frame inputs with a single frame at the question timestamp degenerates model performances, though results remain competitive (which is reasonable as people often ask about their current visual environment.). Figure 5 further suggests that video-level modeling is key for answering questions of “Safety Warning”, “Tool Use”, and “Communication”, while a single frame seems to be sufficient for information reading. Table 4.2 shows that substituting the blind-specific prompt with a normal VQA prompt jeopardizes model performance, highlighting the unique challenges of blind-oriented QA, especially in navigation and tool using as shown in Figure 6. Table 4.2 and Figure 7 demonstrate that using our training data for instruction tuning (LoRA [57] finetune) can remarkably improve performances, yet the gap compared with human remains significant (Implementation details are presented in Appendix Sec. B.3). Additionally, Figure 7 shows that existing finetuning will hurt the performances on “tool use” questions, likely because of overfitting on the limited training data for this category.